# Topology-Driven Parallel Trajectory Optimization in Dynamic Environments

### v1

@@ IV-C Local Planner
(Local Planner) The local planner is an algorithm $L:\mathcal{X}^{N_{L}}\to\mathcal{X}^{N_{L}}$ that respects constraints.
This definition captures many existing optimization-based planners. In this work, we define the local planner through the trajectory optimization in Eq. (III-A), where we make two modifications to ensure that the optimized trajectories are in the homotopy class of the associated guidance trajectory. First, the trajectory optimization of each local planner uses its guidance trajectory as the initial guess for $\bm{x}$. The initial guess speeds up convergence but does not guarantee that the optimized trajectory remains in the same homotopy class when there are obstacles. In the next section (Sec. IV-D), we provide an example where initialization in distinct homotopy classes still leads to identical optimized trajectories.
Figure 6: Topology constraints for each guidance trajectory (blue and green) imposed on the local planner. For a single static obstacle, we visualize for each time step in the trajectory the feasible region (colored regions) and constraints (broken lines) with increasing transparency over time.
@@
Figure 7: Planned trajectories (lines with shaded discs) to avoid a static obstacle with two guidance trajectories (dashed lines) for a low ($0.01$) and high ($0.3$) path folllowing weight. (a) Without homotopy comparison (3), guidance trajectories are not distinct and optimized trajectories are identical. (b) Without homotopy constraints, increasing the path following weight results in identical trajectories. (c) With homotopy comparison (3) and homotopy constraints (11e), optimized trajectories are distinct.
@@
$J_{i}^{*}=\min_{\bm{u}\in\mathbb{U},\bm{x}\in\mathbb{X}}\quad$ $\sum_{k=0}^{N_{L}}J(\bm{x}_{k},\bm{u}_{k})$ (11a)
@@ IV-D Enforcing Consistency over Time
Consider a planning scenario with a robot and static obstacle (both at $y=0$) and the reference path at $y=1$. We plan the robot’s trajectory for a low and high weight on following the path. Fig. 7 shows the planned trajectories after optimization. Without homotopy comparison (3) (see Fig. 6(a)), guidance trajectories are not distinct and lead to identical optimized trajectories. Without homotopy constraints (11e) (see Fig. 6(b)), trajectories are distinct for a low path following weight but become identical by increasing the path following weight. This is possible as the final state of the optimized trajectory is free to move to the other side of the obstacle. Our proposed method (see Fig. 6(c)) maintains the two trajectories in both cases, irrespective of the tuning of the objective function.
@@ IV-E Decision Making
The robot can only execute one trajectory. Since the cost function of the local planners (11a) matches that of the original trajectory optimization (2a), the quality of the guided plans are directly comparable through their optimal costs $J_{i}^{*}$.
@@ IV-F Theoretical Analysis
In the following, we formalize to what extent our proposed planner resolves the nonconvexity of the free space. First, note that due to the cost function and/or nonlinear robot dynamics, the trajectory optimization in Eq. (III-A) remains nonconvex, even when it is constrained to stay in a single homotopy class. There may therefore be multiple local optima in each homotopy class. This means that the proposed planner does not provably return a globally optimal solution to the optimization in Eq. (III-A). We propose instead a weaker notion of globalization [^2] In the following, with some abuse of notation, $J(\bm{\tau})$ refers to the optimal cost of the optimization initialized with trajectory $\bm{\tau}$..
@@
(Homotopy Globally Optimal) Denote the highest-cost local-optimum of optimization (III-A) in homotopy class $i$ as $\bm{\tau}^{-}_{i}$. A trajectory $\bm{\tau}$ is said to be a Homotopy Globally Optimal (HGO) solution if its cost is lower or equal to that solution in each homotopy class, that is, if $J(\bm{\tau})\leq J(\bm{\tau}^{-}_{i}),\ \forall i$, for all homotopy classes that admit a feasible trajectory.
To prove when the proposed scheme computes an HGO solution, we pose three conditions. These conditions link the solution of Eq. (IV-C) to that of Eq. (III-A).
Condition 1. The homotopy constraints are not active ($g_{H}(\bm{x}_{k},\bm{o}_{k}^{j},\bm{\tau}_{i,k})<0\,\ \forall i,k,j$) in the final solution of (IV-C).
@@
If Conditions 1-3 hold, T-MPC is HGO for optimization problem (III-A).
@@
Under Condition $1$, the solution for each optimization (IV-C), $\bm{\tau}_{i}^{*}$, is locally optimal for (III-A) since homotopy constraints (11e) are the only distinction between the two problems. Therefore, $J(\bm{\tau}_{i}^{*})\leq J(\bm{\tau}_{i}^{-})$. If Condition $2$ is satisfied, then $\mathcal{T}_{P}$ contains a guidance trajectory in every feasible homotopy class. Therefore, under Condition $3$, the final trajectory $\bm{\tau}^{*}$ executed by T-MPC satisfies $J(\bm{\tau}^{*})\leq J(\bm{\tau}_{i}^{*})\leq J(\bm{\tau}_{i}^{-}),\ \forall i$ and we obtain the HGO property. ∎
@@ IV-G Non-Guided Local Planner in Parallel
The constraints and initialization provided by the guidance planner allow the local planner to escape poor local optima. Once the planner is in the correct homotopy class, the restrictions imposed by the guidance planner (i.e., homotopy constraints) may degrade performance. For this reason, we consider an extension of the proposed planner where the regular local planner without guidance (i.e., the optimization in Eq. (III-A)) is added to the set of parallel guided local planners. Since this planner is less restricted and does not rely on the global planner, it can occasionally find a better solution.
@@ V-A Implementation
Our implementation for T-MPC is written in C++/ROS and will be released open source. We will also release the guidance planner as a standalone package.
@@
for each time instance $k$ in the horizon $N_{L}$. Herein, $J_{c},J_{l}$ are the contour and lag error used to follow the reference path, $J_{v}=||v-v_{\textrm{ref}}||_{2}^{2}$ tracks a desired velocity and $J_{\omega}=||\omega||_{2}^{2},\ J_{a}=||a||_{2}^{2}$ weigh the control inputs consisting of the rotational velocity $\omega$ and acceleration $a$. Collision avoidance constraints are imposed with $g(\bm{x}_{k},\bm{o}_{k}^{j})\leq 0$,
@@
here $\Delta\bm{p}_{k}^{j}=\bm{p}_{k}-\bm{o}_{k}^{j}$, $\bm{R}(\phi)$ is a rotation matrix with orientation $\phi$ of the robot and $r=r_{\textrm{robot}}+r_{\textrm{obs}}$. These nonconvex constraints directly formulate that the robot region should not overlap with that of the obstacles. We solve each parallel local optimization with Forces Pro [34]. Parameters of the full planner are listed in Table I.
TABLE I: Experimental settings. Weights are denoted “(w)”. Parameter $h$ denotes the discretization time step.
Guidance Planner $N$ $h$ samples $P$ goal grid
$30$ $0.05$ $30$ $4$ $5\times 5$
Local Planner $N$ $h$ $v_{\textrm{ref}}$ $r$
$30$ $0.05$ $2$ m/s $0.725$
Opt. Weights $w_{c}$ $w_{l}$ $w_{v}$ $w_{\omega}$ $w_{a}$
$0.05$ $0.75$ $0.55$ $0.85$ $0.34$
Decision Making Type $c_{i}$
Eq. (14) $0.75$
TABLE II: Quantative results for interactive navigation simulations of Sec. V-C over $200$ experiments with pedestrian motion prediction following a constant velocity model. Task duration (Dur.) and runtime are reported as “mean (std. dev.)”. Best performance in bold.
# Method Dur. [s] Safe (%) Runtime [ms]
4 Motion Primitives [6] 14.0 (0.9) 77 11.6 (2.6)
@@ V-B Simulation Environment
The simulation environment (see Fig. 1) consists of a mobile robot (Clearpath Jackal) moving through a $6$ m wide corridor with up to $12$ pedestrians. The robot follows the centerline with a reference velocity of $2$ m/s and is controlled at $20$ Hz. The pedestrians follow the social forces model [35] using implementation [36]. They interact with other pedestrians and the robot and are aware of the walls. We use a constant velocity model to predict the future pedestrian positions for the planner. The pedestrians have a radius of $0.3$ m. We specify a radius of $0.4$ m in the planners to account for discretization effects, allowing us to clearly identify collisions. Pedestrians spawn on two sides of the corridor with the objective to traverse the corridor. The random start and goal locations are the same for each planner.
@@ V-C Comparison to Baselines
We validate T-MPC and T-MPC++ against four baselines.
  • Motion Primitives (global planner) [6].
  • TEB Local Planner (topology-guided planner) [8]: Using the implementation from the ROS 2D navigation stack [37]. We increased the dynamic object weight from $10$ to $20$ to decrease collisions.
  • LMPCC (local planner) [3]: See Sec. V-A. We supply the previous solution shifted forward in time as the initial guess of the optimization.
  • Guidance-MPCC (topology-guided planner) [28]: We updated the guidance planner to that used in this work to make it more competitive with T-MPC.
@@
Figure 8: Visualization of the task duration (i.e., the time taken to reach the goal) in Table II. Our method achieves the smallest variation in task duration and the shortest task duration in crowded environments.
@@
The results over $200$ experiments are summarized in Table II and the task duration is visualized in Fig. 8. The Motion Primitives planner is not safe and has a much longer task duration than the other planners. LMPCC has a longer task duration than the guided planners and is less safe, in part because LMPCC plans a single trajectory. When the optimization becomes infeasible, it often does not recover fast enough to avoid oncoming obstacles. The TEB Planner outperforms the constrained LMPCC planner, but its safety and task duration degrades quickly when the environment becomes more crowded. Since the TEB planner is soft constrained, it can get pushed into collisions in a crowded environment. Its task duration also varies much more than that of T-MPC and T-MPC++. To further compare T-MPC++ and the TEB Local Planner, we visualize their trajectories in Fig. 9. Our proposed planner results in smoother and more consistent trajectories and can follow the reference path more closely.
@@
Out of the guidance planners, Guidance-MPCC attains the same mean task duration as T-MPC++ but is less consistent (high std. dev.). It often drives faster than the reference velocity since the guidance planner, which does not account for the robot dynamics, determines the planner’s behavior. This also leads to more collisions than LMPCC in all scenarios. T-MPC is generally faster than LMPCC, except for the $12$ pedestrian case. In this environment, the local planner may find solutions that the guidance planner did not, given that the space is cluttered. T-MPC++ demonstrates superior navigation performance over the other planners: it is faster, varies less in its task duration (lower std. dev.) and is safer in almost all cases (the TEB Local Planner is safer in the $8$ pedestrian case) at the cost of a higher runtime. We measured that T-MPC takes $5$ ms on average to compute guidance trajectories in all environments. Finally, T-MPC++ scales better to crowded environments than the baselines.
V-D Sensitivity Studies

### v2

@@ IV-C Local Planner
(Local Planner) The local planner is an algorithm $L:\mathcal{X}^{N}\to\mathcal{X}^{N}$ that respects constraints.
This definition captures many existing optimization-based planners. In this work, we define the local planner through the trajectory optimization in Eq. (2), where we make two modifications to ensure that the optimized trajectories are in the homotopy class of the associated guidance trajectory. First, the trajectory optimization of each local planner uses its guidance trajectory as the initial guess for $\bm{x}$. The initial guess speeds up convergence but does not guarantee that the optimized trajectory remains in the same homotopy class when there are obstacles. In the next section (Sec. IV-D), we provide an example where initialization in distinct homotopy classes still leads to identical optimized trajectories.
(a) Local planner $1$ plans to evade the obstacle left.
(b) Local planner $2$ plans to evade the obstacle right.
Figure 5: Two distinct locally planned trajectories for a robot (black dot) evading an obstacle (black region and dot) that is static ($\bm{o}_{k}=\bm{o}_{0},\ k=1,\ldots,N$). For both planners, we depict the topology constraints for each time step in their respective colors showing the constraint boundaries (broken lines) and their feasible region (colored regions with increasing transparency over time).
@@
Figure 6: Planned trajectories (lines with shaded discs) tracking a reference path (dashed black line) while avoiding a static obstacle with two guidance trajectories (dashed lines) for a low ($0.01$) and high ($0.3$) path following weight. (a) Without homotopy comparison (3), guidance trajectories are not distinct and optimized trajectories are identical. (b) Without homotopy constraints, increasing the path following weight results in identical trajectories. (c) With homotopy comparison (3) and homotopy constraints (9e), optimized trajectories are distinct.
@@
$J_{i}^{*}=\min_{\bm{u}\in\mathbb{U},\bm{x}\in\mathbb{X}}\quad$ $\sum_{k=0}^{N}J(\bm{x}_{k},\bm{u}_{k})$ (9a)
@@ IV-D Enforcing Consistency over Time
Consider a planning scenario with a robot and static obstacle (both at $y=0$) and the reference path at $y=1$. We plan the robot’s trajectory for a low and high weight on following the path [^2] The path following weight is the contouring weight from [3].. Fig. 6 shows the planned trajectories after optimization. Without homotopy comparison (3) (see Fig. 6(a)), guidance trajectories are not distinct and lead to identical optimized trajectories. Without homotopy constraints (9e) (see Fig. 6(b)), trajectories are distinct for a low path following weight but become identical by increasing the path following weight. This is possible as the final state of the optimized trajectory is free to move to the other side of the obstacle. Our proposed method (see Fig. 6(c)) maintains the two trajectories in both cases, irrespective of the tuning of the objective function.
@@ IV-E Decision Making
The robot can only execute one trajectory. Since the cost function of the local planners (9a) matches that of the original trajectory optimization (2a), the quality of the guided plans are directly comparable [^3] As trajectory end points can be distinct, their quality needs to be represented in the cost function. We include a terminal cost that accounts for the deviation of the end point from the reference path. through their optimal costs $J_{i}^{*}$.
@@ IV-F Theoretical Analysis
In the following, we formalize to what extent our proposed planner resolves the nonconvexity of the free space. First, note that due to the cost function and/or nonlinear robot dynamics, the trajectory optimization in Eq. (2) remains nonconvex, even when it is constrained to stay in a single homotopy class. There may therefore be multiple local optima in each homotopy class. This means that the proposed planner does not provably return a globally optimal solution to the optimization in Eq. (2). We propose instead a weaker notion of globalization [^5] In the following, with some abuse of notation, $J(\bm{\tau})$ refers to the optimal cost of the optimization initialized with trajectory $\bm{\tau}$..
@@
(Homotopy Globally Optimal) Denote the highest-cost local-optimum of optimization (2) in homotopy class $i$ as $\bm{\tau}^{-}_{i}$. A trajectory $\bm{\tau}$ is said to be a Homotopy Globally Optimal (HGO) solution if its cost is lower or equal to that solution in each homotopy class, that is, if $J(\bm{\tau})\leq J(\bm{\tau}^{-}_{i}),\ \forall i$, for all homotopy classes that admit a feasible trajectory.
To prove when the proposed scheme computes an HGO solution, we pose three conditions. These conditions link the solution of Eq. (9) to that of Eq. (2).
Condition 1. The homotopy constraints are not active ($g_{H}(\bm{x}_{k},\bm{o}_{k}^{j},\bm{\tau}_{i,k})<0\,\ \forall i,k,j$) in the final solution of (9).
@@
If Conditions 1-3 hold, T-MPC is HGO for optimization problem (2).
@@
Under Condition $1$, the solution for each optimization (9), $\bm{\tau}_{i}^{*}$, is locally optimal for (2) since homotopy constraints (9e) are the only distinction between the two problems. Therefore, $J(\bm{\tau}_{i}^{*})\leq J(\bm{\tau}_{i}^{-})$. If Condition $2$ is satisfied, then $\mathcal{T}_{P}$ contains a guidance trajectory in every feasible homotopy class. Therefore, under Condition $3$, the final trajectory $\bm{\tau}^{*}$ executed by T-MPC satisfies $J(\bm{\tau}^{*})\leq J(\bm{\tau}_{i}^{*})\leq J(\bm{\tau}_{i}^{-}),\ \forall i$ and we obtain the HGO property. ∎
@@ IV-G Non-Guided Local Planner in Parallel
The constraints and initialization provided by the guidance planner allow the local planner to escape poor local optima. Once the planner is in the correct homotopy class, the restrictions imposed by the guidance planner (i.e., homotopy constraints) may degrade performance. For this reason, we consider an extension of the proposed planner where the regular local planner without guidance (i.e., the optimization in Eq. (2)) is added to the set of parallel guided local planners. Since this planner is less restricted and does not rely on the global planner, it can occasionally find a better solution.
@@
IV-H Computation Time Analysis
T-MPC plans guidance trajectories before optimization. In the following, we analyze the computational complexity of the guidance planner (see Algorithm 1), considering the number of PRM samples $n$, obstacles $M$ and distinct trajectories $P$.
Visibility-PRM
The time complexity of regular Visibility-PRM is dominated by the visibility check. When adding a node, it checks its visibility in the worst case to all nodes (if all nodes are Guards), where each visibility check considers all obstacles. Its time complexity therefore is $O(n^{2}M)$. We additionally verify that new connections are distinct. The time complexity of a single homotopy comparison is $O(M)$: the H-signature or winding numbers are evaluated for each obstacle. The homotopy class is compared roughly $n$ times if a new distinct connection neighbors all connectors. Its time complexity therefore is $O(n^{2}M)$ and does not change the time complexity of Visibility-PRM.
DepthFirstSearch
Each node links to at most one goal. Hence searching the graph for at most $P$ paths to each goal at worst considers each node once. Its time complexity is $O(n)$.
FilterAndSelect
Sorting $P$ trajectories has time complexity $O(P\log{P})$. Filtering homotopy distinct trajectories from the sorted list must compare a trajectory to $P$ others in the worst case and has time complexity $O(P^{2}M)$.
IdentifyAndPropagate
Similarly, comparing the homotopy class of new and existing trajectories has time complexity $O(P^{2}M)$. Propagating the graph has time complexity $O(n)$.
Total
The time complexity of the guidance planner is $O((n^{2}+P^{2})M)$. In practice, this time complexity can be approximated by $O(n^{2}M)$, given that the number of relevant homotopy classes is typically small (i.e., $n\gg P$). Thanks to the propagation of nodes from the previous iteration, we find that $n$ can be relatively small as well (e.g., $n<100$). For our use case, a relatively small $n$ and $P$ are usually sufficient to construct a sparse graph from which the relevant homotopy classes can be extracted.
@@ V-A Implementation
Our implementation for T-MPC is written in C++/ROS and will be released open source [^6] See https://github.com/tud-amr/mpc_planner. We will also release the guidance planner as a standalone package.
@@
for each time instance $k$ in the horizon $N$. Herein, $J_{c},J_{l}$ are the contour and lag error used to follow the reference path, $J_{v}=||v-v_{\textrm{ref}}||_{2}^{2}$ tracks a desired velocity and $J_{\omega}=||\omega||_{2}^{2},\ J_{a}=||a||_{2}^{2}$ weigh the control inputs consisting of the rotational velocity $\omega$ and acceleration $a$. Collision avoidance constraints are imposed with $g(\bm{x}_{k},\bm{o}_{k}^{j})\leq 0$,
@@
here $\Delta\bm{p}_{k}^{j}=\bm{p}_{k}-\bm{o}_{k}^{j}$, $\bm{R}(\phi)$ is a rotation matrix with orientation $\phi$ of the robot and $r=r_{\textrm{robot}}+r_{\textrm{obs}}$. These nonconvex constraints directly formulate that the robot region should not overlap with that of the obstacles. We solve each parallel local optimization with Forces Pro [41]. Parameters of the full planner are listed in Table I. Weights of the guidance and local planners are manually tuned. The planning scheme, including guidance and local planners, is updated in each iteration in a receding horizon manner.
TABLE I: Experimental settings.
Parameter Name Parameter Value Parameter Description
$N$ $30$ Global and local planner horizon
$\Delta T$ $0.2$ s Integration time step
$h$ $0.05$ s Planning time step
  $n$ $30$ Visibility-PRM sample limit
$T_{\textrm{max}}$ $10$ ms Visibility-PRM time limit
Eq. 3 H-signature Homotopy comparison function
$P$ $4$ of distinct guidance trajectories
$G$ $5\times 5$ Grid of goals (longitudinal $\times$ lateral)
  $r$ $0.725$ m Combined obstacle and robot radius
$w_{c}$ $0.05$ Optimization contouring weight
$w_{l}$ $0.75$ Optimization lag weight
$w_{v}$ $0.55$ Optimization velocity tracking weight
$w_{\omega}$ $0.85$ Optimization rotational velocity weight
$w_{a}$ $0.34$ Optimization acceleration weight
  Decision Eq. (12) Type of decision-making
$c_{i}$ $0.75$ Discount factor for trajectory in previously followed homotopy class
TABLE II: Quantative results for interactive navigation simulations of Sec. V-C over $200$ experiments with pedestrian motion prediction following a constant velocity model. Task duration (Dur.) and runtime are reported as “mean (std. dev.)”. Without obstacles, the task duration is $12.9$ s. Best planner performances per column are denoted in bold. Underlined results indicate that T-MPC++ significantly outperforms the respective method as tested with a U-test for a significance value of $p=0.001$.
# Ped. Method Dur. [s] Safe (%) Runtime [ms]
00 - 12.9 (0.0) - -
4 Frenét-Planner [6] 14.0 (0.9) 77 11.6 (2.6)
@@ V-B Simulation Environment
The first simulation environment (see Fig. 1) consists of a mobile robot (Clearpath Jackal) moving through a $6$ m wide corridor with up to $12$ pedestrians. The robot follows the centerline with a reference velocity of $2$ m/s and is controlled at $20$ Hz. The pedestrians follow the social forces model [42] using implementation [43]. They interact with other pedestrians and the robot and are aware of the walls. We use a constant velocity model to predict the future pedestrian positions for the planner. The pedestrians have a radius of $0.3$ m. We specify a radius of $0.4$ m in the planners to account for discretization effects, allowing us to clearly identify collisions. Pedestrians spawn on two sides of the corridor with the objective to traverse the corridor. The random start and goal locations are the same for each planner.
@@ V-C Comparison to Baselines
We compare T-MPC and T-MPC++ against four baselines. Baselines are selected on the availability of an open-source implementation and their application to navigation in $2$D dynamic environments. We consider the following baselines:
  • Motion Primitives (global planner) [6]: A non-optimization-based global planner that respects the robot dynamics.
  • TEB Local Planner (topology-guided planner) [8]: One of the most used local planners in the ROS navigation stack [44] that considers multiple homotopy classes.
  • LMPCC (local planner) [3]: An open-source non parallelized MPC (see Sec. V-A). We supply the previous solution shifted forward in time as the initial guess of the optimization.
  • Guidance-MPCC (topology-guided planner) [32]: Our previous conference work. We updated the guidance planner to that used in this work to make it more competitive with T-MPC++.
We use the same weights for the MPC planners (LMPCC, Guidance-MPCC, T-MPC, T-MPC++). Baseline actuation limits and tracking objectives are adapted to match the MPC objectives. TEB Local Planner tuning uses its default but with increased collision avoidance weight (from $10$ to $20$) to decrease collisions in crowded environments.
@@
Figure 7: Visualization of the task duration (i.e., the time taken to reach the goal) in Table II. The dashed vertical line denotes the task duration without obstacles. Our method achieves the smallest variation in task duration and the shortest task duration in crowded environments.
@@
The results over $200$ experiments are summarized in Table II and the task duration is visualized in Fig. 7. The Motion Primitives planner is not safe and has a significantly longer task duration than the other planners even in the least crowded case. In the two more crowded environments, LMPCC has a significantly longer task duration than T-MPC++ and is less safe, in part because LMPCC plans a single trajectory. When the optimization becomes infeasible, it often does not recover fast enough to avoid oncoming obstacles. TEB Local Planner is marginally safer than LMPCC and maintains competitive average task durations to the other methods in the four and eight pedestrian scenarios under less computational cost. In the crowded scenario, T-MPC++ completes the task significantly faster. Additionally, in all cases, T-MPC++ has a much smaller standard deviation of the task duration indicating that its behavior is more consistent (visible also in Fig. 7). The TEB Local Planner soft constrains collision avoidance which, in crowded environments, leads the robot into poor behaviors (e.g., reversing) due to the shape of the cost function. To further compare T-MPC++ and the TEB Local Planner, we visualize their trajectories in Fig. 8. Our proposed planner results in smoother and more consistent trajectories and can follow the reference path more closely. We quantitatively compare the smoothness of the planners through the standard deviation on second-order input commands. The deviation on acceleration ($\sigma_{a}$) and rotational acceleration ($\sigma_{\alpha}$) for TEB Local Planner are higher ($\sigma_{a}=0.16$, $\sigma_{\alpha}=0.12$) than for T-MPC++ ($\sigma_{a}=0.04$, $\sigma_{\alpha}=0.05$).
@@
Out of the guidance planners, Guidance-MPCC attains the same mean task duration as T-MPC++ but is less consistent (high std. dev.) because the guidance planner, which does not account for the robot dynamics, determines the planner’s behavior. This results in larger tracking errors under the same cost function, for example in the $12$ pedestrian case, the mean path and velocity errors of Guidance-MPCC are $0.45$m and $0.47$m/s, respectively, compared to values of $0.20$m and $0.42$m/s for T-MPC++. It also leads to more collisions than LMPCC. T-MPC is generally faster than LMPCC, except for the $12$ pedestrian case. In this environment, the local planner may find solutions that the guidance planner did not, given that the space is cluttered. T-MPC++ demonstrates superior navigation performance over the other planners: it is significantly faster (with the exception of Guidance-MPCC), varies less in its task duration (lower std. dev.) and is safer in almost all cases (the TEB Local Planner is safer in the $8$ pedestrian case). T-MPC++ has higher computational demands than the other planners. Compared to the other MPC planners, T-MPC++ first computes guidance plans. We measured that this step takes approximately $5$ ms on average (included in the runtime of Table II) in all scenarios.
V-D Crowded Baseline Comparison
We further compare T-MPC++ against TEB Local Planner in a square-shaped crowded environment with $50$ pedestrians. The robot’s task is to move diagonally through the environment at $v_{\textrm{ref}}=1.5$ m/s. We consider the $12$ pedestrians closest to the plan, with preference for nearby pedestrians in both methods. Pedestrians are removed when they reach the goal to prevent unpredictable turns. Table III presents the results. T-MPC++ is significantly faster and the standard deviation of the task duration is less than half that of TEB Local Planner. TEB Local Planner is on average computationally faster as it does not compute a new plan in each iteration. When it does compute a plan, its computation time can exceed the planning frequency of $20$ Hz as shown by the maximum in Table III (it exceeded $50$ ms in $0.26$% of its iterations). In contrast, T-MPC++ is explicitly limited to $50$ ms such that its maximum computation time remains below $50$ ms.
TABLE III: Quantative results in crowded environment of Sec. V-D over $200$ experiments. Notation follows that of Table II. Without obstacles, the task duration is $20.6$ s. The runtime is denoted as “mean (max)”.
# Ped. Method Dur. [s] Safe (%) Runtime [ms]
00 - 20.6 (0.0) - -
50 TEB Local Planner [8] 22.4 (2.4) 92 9.4 (177.1)
T-MPC++ (ours) 21.0 (0.9) 92 21.2 (46.9)
V-E Sensitivity Studies